# Large Language Models Are Zero-Shot Time Series Forecasters

### v1

@@ 1 Introduction
Despite similarities with other sequence modeling problems, such as text, audio, or video, time series has two particularly challenging properties. Unlike video or audio, which typically have consistent input scales and sampling rates, aggregated time series datasets often comprise sequences from radically different sources, sometimes with missing values. Moreover, common applications of time series forecasting, such as weather or financial data, require extrapolating from observations that contain a tiny fraction of the possible information, making accurate point predictions nearly impossible and uncertainty estimation especially important. While large-scale pretraining has become a key element of training large neural networks in vision and text, enabling performance to scale directly with data availability, pretraining is not typically used for time series modeling, where there is no consensus unsupervised objective and large, cohesive pretraining datasets are not readily available. Consequently, simple time series methods (e.g. ARIMA (Box and Jenkins, 1968), and linear models (Zeng et al., 2022)) often outperform deep learning methods on popular benchmarks (Hewamalage et al., 2023).
@@
To understand the origins of LLMTime’s impressive performance, we investigate how LLMs express preferences for simple or repetitive sequences (Goldblum et al., 2023) and show that these biases are in fact compatible with the salient structure of time series, such as seasonality. Aside from these biases, LLMs also can naturally acccommodate missing data, and express multimodal distributions, which is particularly useful for time series. We also show how LLMs enable appealing functionality, such as the ability to provide additional side information through prompting, and query the LLM to explain its predictions.
@@ 2 Background
Language modeling    Language models are trained on a collection of sequences, $\mathcal{U}={\{U_{1},U_{2},\dots U_{i},\dots,U_{N}\}}$, where $U_{i}=(u_{1},u_{2},\dots,u_{j},\dots,u_{n_{i}})$ and each token, $u_{i}$, belongs to a vocabulary $\mathcal{V}$. Large language models typically encode an autoregressive distribution, in which the probability of each token is only dependent on the previous tokens in the sequence, $p_{\theta}\left(U_{i}\right)=\prod_{j=1}^{n_{i}}p_{\theta}\left(u_{j}\mid u_{0:j-1}\right)$. The parameters, $\theta$, are learned by maximizing the probability of the entire dataset ${p_{\theta}(\mathcal{U})=\prod_{i=1}^{N}p_{\theta}(U_{i})}$. Every language model has an associated tokenizer, which breaks an input string into a sequence of tokens, each belonging to $\mathcal{V}$. Proper tokenization is extremely important, and small details can have surprisingly significant effects. The most common tokenization method for autoregressive language models is byte-pair encoding (BPE), which treats inputs like bit strings and assigns tokens based on the rate of occurrence in the training corpus, optimizing for shorter sequences of tokens on average. Sampling from a language model typically starts with a prompt, $u_{0:k}$, and proceeds sequentially using $p_{\theta}\left(u_{j}\mid u_{0:j-1}\right)$, which is often preprocessed, for example through temperature scaling or nucleus sampling (Holtzman et al., 2019).
@@
Large language models    Brown et al. (2020) showed that increasing a language model’s parameter count and training data size leads to new capabilities such as zero-shot generalization, in which a model can perform a text-formatted task without training the model parameters on any task-specific data. Large language models, for example GPT-3 (Brown et al., 2020) or LLaMA-2 (Touvron et al., 2023b), accomplish this form of generalization through in-context learning, which identifies patterns in the language model’s prompt and extrapolates them through next-token prediction. Many authors have speculated that in-context learning emerges from a language model’s extensive compression of the input data (Goldblum et al., 2023; Sutskever, 2023; Delétang et al., 2023). Compression favors learning algorithms that operate over the input data with programmatic abstractions, for example, context-free grammars (Allen-Zhu and Li, 2023) or induction heads (Olsson et al., 2022), which can implement copy-and-paste type operations for generating samples with highly structured syntax. In this work, we show that the zero-shot generalization abilities of LLMs and their preference for compressible patterns extend well beyond language understanding and can be used for time series forecasting.
@@
Zero-shot generalization has made LLMs significantly more useful as assistants, leading to the create of methods to align LLMs with human preferences and instructions, for example reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) and instruction tuning (Wei et al., 2021). While key to modern LLMs products, alignment methods can also significantly affect the abilities and calibration of the underlying model (OpenAI, 2023; Bubeck et al., 2023). Here we show these methods can also affect forecasting ability.
@@
Time series data    Time series data typically takes the exact same form as language modeling data, as a collection of sequences $\{U_{i}=(u_{1},\dots,u_{j},\dots,u_{n_{i}})\}$, but in time series $u_{j}$ is numerical. Because language models are built to represent complex probability distributions over sequences, they are theoretically well-suited for time series modeling. In practice, however, language models are held back by the details of tokenizing numbers. BPE compresses numbers based on frequency of occurrence in the training data, so numbers can be broken down into awkward chunks that make learning basic numerical operations challenging. Touvron et al. (2023a) therefore designed the LLaMA tokenizer to map numbers to individual digits, which can lead to significant improvements in mathematical abilities, with small LLaMA models outperforming GPT-4 (Liu and Low, 2023).
@@
Language models for time series    Several authors have explored using pretrained language model encoders as initializations for time series models. For example, Zhou et al. (2023) propose FPT, which finetunes a BERT encoder to perform time series forecasting. Similarly, Zhang et al. (2023) introduce Meta-Transformer, a framework for finetuning a language model for non-text modalities, including time series. Fewer papers explore using LLMs as forecasters without finetuning. The only method we are aware of is PromptCast (Xue and Salim, 2023), which poses forecasting as question answering with prompting.
@@ 3 LLMTime: Forecasting with Language Models
Tokenization    Tokenization is particularly important because it directly influences how patterns form within tokenized sequences and the types of operations that language models can learn. Unfortunately, common tokenization methods like BPE tend to break a single number into tokens that don’t align with the digits, which can make arithmetic considerably more difficult (Liu and Low, 2023). For example, the number $42235630$ gets tokenized as $[422,35,630]$ by the GPT-3 tokenizer, and changes by even a single digit can result in an entirely different tokenization. By contrast, in many new open-source LLMs (e.g. LLaMA (Touvron et al., 2023a)), numbers are tokenized into individual digits by default. To remedy the tokenization of GPT models, we separate the digits with spaces to force a separate tokenization of each digit and use a comma (" ,") to separate each time step in a time series. Because decimal points are redundant given a fixed precision, we drop them in the encoding to save on context length. Thus, with e.g. $2$ digits of precision, we pre-process a time series as follows before feeding into the tokenizer:
@@
Figure 2: Careful tokenization is important for good forecasting with LLMs. Using the Australian Wine dataset from Darts (Herzen et al., 2022), with values [151, 167, …, 267], we show the tokenization used by GPT-3 (Brown et al., 2020) and LLaMA-2 (Touvron et al., 2023b) and the corresponding effect on forecasting performance. Added spaces allow GPT-3 to create one token per digit, leading to good performance. LLaMA-2, on the other hand, tokenizes digits individually, and adding spaces hurts performance.
@@
Continuous likelihoods    Modeling sequences of individual digits has additional benefits beyond good samples. With $n$ digits of precision in base $B$, each sequence of $n$ digits after the decimal place corresponds to one of $B^{n}$ possible bins (Figure 3), each with width $B^{-n}$. As each distribution $p(u_{j}\mid u_{0:j-1};\theta)$ is a softmax over possible digits, we can view the distribution over each individual number as a hierarchical softmax (Mnih and Hinton, 2008), with $p(u_{1},...,u_{n})=p(u_{n}|u_{n-1},...,u_{0})\,p(u_{1}|u_{0})\,p(u_{0})$. Though a language model’s probability distribution is discrete, we can easily adapt it to provide a continuous density by placing a uniform distribution in each bin. Enumerating each of the countably infinite numbers that the model can produce (because the model can output an arbitrary number of digits before the decimal point) with an index $k\in\mathbb{N}$ each with probability $p_{k}$, we can write out the distribution as a mixture of disjoint uniform distributions over the bins $p(x)=\sum_{k\in\mathbb{N}}p_{k}U_{k}(x)$ where $U_{k}(x)=B^{n}\mathbb{I}_{x\in[B^{-n}k,B^{-n}(k+1))}$. Therefore if a given data point lies in bin $k$, its continuous log likelihood is $\log p(x)=\log p_{k}+n\log B$. Finally, to obtain the likelihood $\log p(z)$ in the original input space, we add a change of variables factor $\log|\tfrac{dx}{dz}|,$ where $z\mapsto x=s(z)$ is the rescaling operation in the pre-processing. As a result, the exponentially large number of bins and exponentially small bin widths enabled by our construction make it surprisingly efficient to represent flexible and high-resolution continuous distributions with LLMs, despite using a discrete tokenization of numbers.
@@
Language models as flexible distributions    The fact that LLMs can express flexible distributions over numbers is key for time series data. Uncertainty quantification is essential to forecasting, and typical approaches to representing uncertainty in time series can be limited by misspecification. For example, one common method for creating a probabilistic forecast is to fit a Gaussian or Laplace observation model. When the underlying data distribution is multimodal, both of these models will perform poorly. Methods like Gaussian mixture models (GMMs) solve the issue of multimodality but introduce additional challenges to optimization and model selection. We show that a language model is an underrated solution by training a small autoregressive model on a variety of one-dimensional distributions shown in Figure 3 (right). These distributions come from an exponential random variable, a mixture of a uniform and a student-t distribution, and the heavy-tailed distribution of time series prediction residuals from an ARIMA model on the MonthlyMilk dataset (Herzen et al., 2022). We evaluate these fits quantitatively by computing Wasserstein distances, and compare to a Laplace observation model, a GMM trained with expectation-maximization, and logistic regression over a flat binning of the data (with a tuned bin size). Each model is trained with only $200$ samples from the distribution. The results show that the decimal autoregressive language model (“Decimal AR”) performs extremely well, handling asymmetric, multimodal, and heavy-tailed distributions, which are among the diverse types characteristic of time series data.
@@
Figure 3: (left): Autoregressive models over sequences of digits act like hierarchical softmax distributions over the corresponding numbers. When combined with uniform distributions in each discrete bucket, distributions over strings can become expressive distributions over continuous domains (right): Using simple autoregressive models (e.g. RNNs) trained on a string representation of numbers, we can fit complex distributions that can be challenging for other methods, such as heavy-tailed or multimodal distributions. A simple autoregressive model can match or outperform well-known methods for density estimation, such as Gaussian mixture models (GMMs) or binning with a fixed resolution, as measured by Wasserstein distance between samples.
@@ 4 Experiments
  • Darts (Herzen et al., 2022): A collection of 8 real univariate time series datasets. For Darts, we use several methods that are implemented directly in the package, including neural network models (TCN (Lea et al., 2016), N-BEATS (Oreshkin et al., 2020), N-HiTS (Challu et al., 2022)) and simple moving average models (ARIMA (Box and Jenkins, 1968)). Darts enables learning observation models with tractable likelihoods and is therefore especially useful for benchmarking the probabilistic predictions of LLMTime. We also include Spectral Mixture Gaussian Process (SM-GP) (Wilson and Adams, 2013), a Bayesian nonparametric method (details in Appendix C.1).
  • Monash (Godahewa et al., 2021): The Monash forecasting archive contains 30 publicly available datasets along with baseline numbers for 12 forecasting models, including simple exponential smooth (e.g. ETS (Hyndman et al., 2008)), gradient boosting (e.g. CatBoost (Prokhorenkova et al., 2018)) and deep learning models (e.g. DeepAR (Salinas et al., 2020), WaveNet (Oord et al., 2016)). The Monash archive comprises over 400,000 individual time series, making it infeasible to use in its entirety with the largest available LLMs. To reduce the computational burden, we evaluate GPT-3’s zero-shot performance on the top 16 datasets where the baseline models perform the best (details in Appendix C.2).
  • Informer (Zhou et al., 2021): We evaluated on multivariate datasets widely used for benchmarking efficient transformer models (Du et al., 2022; Zhou et al., 2021). In order to predict multivariate data with LLMTime, we forecast each covariate independently. We baseline against numbers obtained by running public implementations from the Autoformer (Wu et al., 2021) and FEDFormer (Zhou et al., 2022) codebases (Appendix C.3).
Figure 4: LLMTime with base model GPT-3 or LLaMA-2 70B has the best or second best aggregated performance on several deterministic time series benchmarks (Herzen et al., 2022; Godahewa et al., 2021; Zhou et al., 2021) while being entirely zero-shot. Collectively, these benchmarks comprise 29 individual datasets with diverse sources, lengths, and noise levels. For Monash MAE numbers, established results are reported on unnormalized data, so we normalize values before aggregating (Appendix C.2). The informer datasets are multivariate, and we predict each covariate independently with LLMTime (Appendix C.3). GPT-3 evaluation on the Informer datasets was skipped because of the cost of API queries. Error bars show standard errors over the individual datasets in each benchmark.
@@
Comparison with PromptCast    Though included in the results described above, we want to explicitly highlight that LLMTime significantly outperforms PromptCast (Xue and Salim, 2023) when applied to both GPT-3 and LLaMA-2 70B, according to CRPS and MAE aggregated over the Darts datasets. This performance gap highlights important differences between the two approaches. Unlike our method, PromptCast formulates forecasting as a conventional question-answering task in NLP by prompting pre-trained language models with an explicit question about future values in a time series. For example, PromptCast feeds in the prompt “The values in the WoolyDataset for the past 95 time steps are 6172, 6709, 6633, … , 6077. What will the values for the next 5 time steps be? The values for the next 5 time steps will be", to extract predictions from an LLM. PromptCast also does not apply our tokenization and data rescaling strategy (Section 3), which we show is crucial for good performance.
Figure 5: Extended experiments on the Darts datasets. (left): Example probabilistic forecasts with baseline negative log likelihood per dimension (NLL/D). LLMs easily extrapolate trends (e.g. AirPassengers) and reproduce local patterns when data is noisy (e.g. GasRateCO2). (center): When using probabilistic metrics like NLL and CRPS, LLMTime outperforms all baselines, including PromptCast (Xue and Salim, 2023), a competing LLM method. Error bars show standard errors over datasets with Darts. (right): LLMTime is much more sample efficient than competing methods. While the performance of other methods degrades rapidly when we restrict them to a fraction of the original training set, LLMTime can assign high likelihood with only a few examples.

### v2

@@ 1 Introduction
Despite similarities with other sequence modeling problems, such as text, audio, or video, time series has two particularly challenging properties. Unlike video or audio, which typically have consistent input scales and sampling rates, aggregated time series datasets often comprise sequences from radically different sources, sometimes with missing values. Moreover, common applications of time series forecasting, such as weather or financial data, require extrapolating from observations that contain a tiny fraction of the possible information, making accurate point predictions nearly impossible and uncertainty estimation especially important. While large-scale pretraining has become a key element of training large neural networks in vision and text, enabling performance to scale directly with data availability, pretraining is not typically used for time series modeling, where there is no consensus unsupervised objective and large, cohesive pretraining datasets are not readily available. Consequently, simple time series methods (e.g. ARIMA [8], and linear models [52]) often outperform deep learning methods on popular benchmarks [24].
@@
To understand the origins of LLMTime’s impressive performance, we investigate how LLMs express preferences for simple or repetitive sequences [19] and show that these biases are in fact compatible with the salient structure of time series, such as seasonality. Aside from these biases, LLMs also can naturally acccommodate missing data, and express multimodal distributions, which is particularly useful for time series. We also show how LLMs enable appealing functionality, such as the ability to provide additional side information through prompting, and query the LLM to explain its predictions.
@@ 2 Background
Language modeling    Language models are trained on a collection of sequences, $\mathcal{U}={\{U_{1},U_{2},\dots U_{i},\dots,U_{N}\}}$, where $U_{i}=(u_{1},u_{2},\dots,u_{j},\dots,u_{n_{i}})$ and each token, $u_{i}$, belongs to a vocabulary $\mathcal{V}$. Large language models typically encode an autoregressive distribution, in which the probability of each token is only dependent on the previous tokens in the sequence, $p_{\theta}\left(U_{i}\right)=\prod_{j=1}^{n_{i}}p_{\theta}\left(u_{j}\mid u_{0:j-1}\right)$. The parameters, $\theta$, are learned by maximizing the probability of the entire dataset ${p_{\theta}(\mathcal{U})=\prod_{i=1}^{N}p_{\theta}(U_{i})}$. Every language model has an associated tokenizer, which breaks an input string into a sequence of tokens, each belonging to $\mathcal{V}$. Proper tokenization is extremely important, and small details can have surprisingly significant effects. The most common tokenization method for autoregressive language models is byte-pair encoding (BPE), which treats inputs like bit strings and assigns tokens based on the rate of occurrence in the training corpus, optimizing for shorter sequences of tokens on average. Sampling from a language model typically starts with a prompt, $u_{0:k}$, and proceeds sequentially using $p_{\theta}\left(u_{j}\mid u_{0:j-1}\right)$, which is often preprocessed, for example through temperature scaling or nucleus sampling [25].
@@
Large language models    Brown et al. [9] showed that increasing a language model’s parameter count and training data size leads to new capabilities such as zero-shot generalization, in which a model can perform a text-formatted task without training the model parameters on any task-specific data. Large language models, for example GPT-3 [9] or LLaMA-2 [44], accomplish this form of generalization through in-context learning, which identifies patterns in the language model’s prompt and extrapolates them through next-token prediction. Many authors have speculated that in-context learning emerges from a language model’s extensive compression of the input data [19, 42, 15]. Compression favors learning algorithms that operate over the input data with programmatic abstractions, for example, context-free grammars [2] or induction heads [34], which can implement copy-and-paste type operations for generating samples with highly structured syntax. In this work, we show that the zero-shot generalization abilities of LLMs and their preference for compressible patterns extend well beyond language understanding and can be used for time series forecasting.
@@
Zero-shot generalization has made LLMs significantly more useful as assistants, leading to the create of methods to align LLMs with human preferences and instructions, for example reinforcement learning from human feedback (RLHF) [38] and instruction tuning [46]. While key to modern LLMs products, alignment methods can also significantly affect the abilities and calibration of the underlying model [36, 10]. Here we show these methods can also affect forecasting ability.
@@
Time series data    Time series data typically takes the exact same form as language modeling data, as a collection of sequences $\{U_{i}=(u_{1},\dots,u_{j},\dots,u_{n_{i}})\}$, but in time series $u_{j}$ is numerical. Because language models are built to represent complex probability distributions over sequences, they are theoretically well-suited for time series modeling. In practice, however, language models are held back by the details of tokenizing numbers. BPE compresses numbers based on frequency of occurrence in the training data, so numbers can be broken down into awkward chunks that make learning basic numerical operations challenging. Touvron et al. [43] therefore designed the LLaMA tokenizer to map numbers to individual digits, which can lead to significant improvements in mathematical abilities, with small LLaMA models outperforming GPT-4 [30].
@@
Language models for time series    Several authors have explored using pretrained language model encoders as initializations for time series models. For example, Zhou et al. [56] propose FPT, which finetunes a BERT encoder to perform time series forecasting. Similarly, Zhang et al. [53] introduce Meta-Transformer, a framework for finetuning a language model for non-text modalities, including time series. Fewer papers explore using LLMs as forecasters without finetuning. The only method we are aware of is PromptCast [50], which poses forecasting as question answering with prompting.
@@ 3 LLMTime: Forecasting with Language Models
Tokenization    Tokenization is particularly important because it directly influences how patterns form within tokenized sequences and the types of operations that language models can learn. Unfortunately, common tokenization methods like BPE tend to break a single number into tokens that don’t align with the digits, which can make arithmetic considerably more difficult [30]. For example, the number $42235630$ gets tokenized as $[422,35,630]$ by the GPT-3 tokenizer, and changes by even a single digit can result in an entirely different tokenization. By contrast, in many new open-source LLMs (e.g. LLaMA [43]), numbers are tokenized into individual digits by default. To remedy the tokenization of GPT models, we separate the digits with spaces to force a separate tokenization of each digit and use a comma (" ,") to separate each time step in a time series. Because decimal points are redundant given a fixed precision, we drop them in the encoding to save on context length. Thus, with e.g. $2$ digits of precision, we pre-process a time series as follows before feeding into the tokenizer:
@@
Figure 2: Careful tokenization is important for good forecasting with LLMs. Using the Australian Wine dataset from Darts [23], with values [151, 167, …, 267], we show the tokenization used by GPT-3 [9] and LLaMA-2 [44] and the corresponding effect on forecasting performance. Added spaces allow GPT-3 to create one token per digit, leading to good performance. LLaMA-2, on the other hand, tokenizes digits individually, and adding spaces hurts performance.
@@
Continuous likelihoods    Modeling sequences of individual digits has additional benefits beyond good samples. With $n$ digits of precision in base $B$, each sequence of $n$ digits after the decimal place corresponds to one of $B^{n}$ possible bins (Figure 3), each with width $B^{-n}$. As each distribution $p(u_{j}\mid u_{0:j-1};\theta)$ is a softmax over possible digits, we can view the distribution over each individual number as a hierarchical softmax [31], with $p(u_{1},...,u_{n})=p(u_{n}|u_{n-1},...,u_{0})\,p(u_{1}|u_{0})\,p(u_{0})$. Though a language model’s probability distribution is discrete, we can easily adapt it to provide a continuous density by placing a uniform distribution in each bin. Enumerating each of the countably infinite numbers that the model can produce (because the model can output an arbitrary number of digits before the decimal point) with an index $k\in\mathbb{N}$ each with probability $p_{k}$, we can write out the distribution as a mixture of disjoint uniform distributions over the bins $p(x)=\sum_{k\in\mathbb{N}}p_{k}U_{k}(x)$ where $U_{k}(x)=B^{n}\mathbb{I}_{x\in[B^{-n}k,B^{-n}(k+1))}$. Therefore if a given data point lies in bin $k$, its continuous log likelihood is $\log p(x)=\log p_{k}+n\log B$. Finally, to obtain the likelihood $\log p(z)$ in the original input space, we add a change of variables factor $\log|\tfrac{dx}{dz}|,$ where $z\mapsto x=s(z)$ is the rescaling operation in the pre-processing. As a result, the exponentially large number of bins and exponentially small bin widths enabled by our construction make it surprisingly efficient to represent flexible and high-resolution continuous distributions with LLMs, despite using a discrete tokenization of numbers.
@@
Language models as flexible distributions    The fact that LLMs can express flexible distributions over numbers is key for time series data. Uncertainty quantification is essential to forecasting, and typical approaches to representing uncertainty in time series can be limited by misspecification. For example, one common method for creating a probabilistic forecast is to fit a Gaussian or Laplace observation model. When the underlying data distribution is multimodal, both of these models will perform poorly. Methods like Gaussian mixture models (GMMs) solve the issue of multimodality but introduce additional challenges to optimization and model selection. We show that a language model is an underrated solution by training a small autoregressive model on a variety of one-dimensional distributions shown in Figure 3 (right). These distributions come from an exponential random variable, a mixture of a uniform and a student-t distribution, and the heavy-tailed distribution of time series prediction residuals from an ARIMA model on the MonthlyMilk dataset [23]. We evaluate these fits quantitatively by computing Wasserstein distances, and compare to a Laplace observation model, a GMM trained with expectation-maximization, and logistic regression over a flat binning of the data (with a tuned bin size). Each model is trained with only $200$ samples from the distribution. The results show that the decimal autoregressive language model (“Decimal AR”) performs extremely well, handling asymmetric, multimodal, and heavy-tailed distributions, which are among the diverse types characteristic of time series data.
@@
Figure 3: Left: Autoregressive models over sequences of digits act like hierarchical softmax distributions over the corresponding numbers. When combined with uniform distributions in each discrete bucket, distributions over strings can become expressive distributions over continuous domains. Right: Using simple autoregressive models (e.g. RNNs) trained on a string representation of numbers, we can fit complex distributions that can be challenging for other methods, such as heavy-tailed or multimodal distributions. A simple autoregressive model can match or outperform well-known methods for density estimation, such as Gaussian mixture models (GMMs) or binning with a fixed resolution, as measured by Wasserstein distance between samples.
@@ 4 Experiments
  • Darts [23]: A collection of 8 real univariate time series datasets. For Darts, we use several methods that are implemented directly in the package, including neural network models (TCN [28], N-BEATS [37], N-HiTS [11]) and simple moving average models (ARIMA [8]). Darts enables learning observation models with tractable likelihoods and is therefore especially useful for benchmarking the probabilistic predictions of LLMTime. We also include Spectral Mixture Gaussian Process (SM-GP) [48], a Bayesian nonparametric method (details in Appendix C.1).
  • Monash [18]: The Monash forecasting archive contains 30 publicly available datasets along with baseline numbers for 12 forecasting models, including simple exponential smooth (e.g. ETS [26]), gradient boosting (e.g. CatBoost [39]) and deep learning models (e.g. DeepAR [40], WaveNet [35]). The Monash archive comprises over 400,000 individual time series, making it infeasible to use in its entirety with the largest available LLMs. To reduce the computational burden, we evaluate GPT-3’s zero-shot performance on 19 datasets described in Appendix C.2.
  • Informer [54]: We evaluated on multivariate datasets widely used for benchmarking efficient transformer models [16, 54]. In order to predict multivariate data with LLMTime, we forecast each covariate independently. We baseline against numbers obtained by running public implementations from the Autoformer [49] and FEDFormer [55] codebases (Appendix C.3).
Figure 4: LLMTime with base model GPT-3 or LLaMA-2 70B has the best or second best aggregated performance on several deterministic time series benchmarks [23, 18, 54] while being entirely zero-shot. Collectively, these benchmarks comprise 29 individual datasets with diverse sources, lengths, and noise levels. For Monash MAE numbers, established results are reported on unnormalized data, so we normalize values before aggregating (Appendix C.2). The informer datasets are multivariate, and we predict each covariate independently with LLMTime (Appendix C.3). GPT-3 evaluation on the Informer datasets was skipped because of the cost of API queries. Error bars show standard errors over the individual datasets in each benchmark.
@@
Comparison with PromptCast    Though included in the results described above, we want to explicitly highlight that LLMTime significantly outperforms PromptCast [50] when applied to both GPT-3 and LLaMA-2 70B, according to CRPS and MAE aggregated over the Darts datasets. This performance gap highlights important differences between the two approaches. Unlike our method, PromptCast formulates forecasting as a conventional question-answering task in NLP by prompting pre-trained language models with an explicit question about future values in a time series. For example, PromptCast feeds in the prompt “The values in the WoolyDataset for the past 95 time steps are 6172, 6709, 6633, … , 6077. What will the values for the next 5 time steps be? The values for the next 5 time steps will be", to extract predictions from an LLM. PromptCast also does not apply our tokenization and data rescaling strategy (Section 3), which we show is crucial for good performance.
Figure 5: Extended experiments on the Darts datasets. Left: Example probabilistic forecasts with baseline negative log likelihood per dimension (NLL/D). LLMs easily extrapolate trends (e.g. AirPassengers) and reproduce local patterns when data is noisy (e.g. GasRateCO2). Center: When using probabilistic metrics like NLL and CRPS, LLMTime outperforms all baselines, including PromptCast [50], a competing LLM method. Error bars show standard errors over datasets with Darts. Right: LLMTime is much more sample efficient than competing methods. While the performance of other methods degrades rapidly when we restrict them to a fraction of the original training set, LLMTime can assign high likelihood with only a few examples.